Assume an income equation, $24 = $6 Meat + $3 Cheese, and an appropriate indifference curve. Solve the intercept points and graph my budget constraint for these two goods. If you did Entry 10, you may use those numbers.

Let the price of Meat falls from $6 to $5 and then to $4.

Draw the resulting budget constraints (with indifference curves) and derive the Demand Curve for Meat.
Answer #1

Let Meat be on (Horizontal) X axis and Cheese be on (Vertical) Y axis.

X axis intercept= Income/Price of meat= 24/6=4

Y axis intercept= Income/ price of cheese= 24/3=8

Intercepts are (4,0)&(0,8).

Budget line BL1= 24=6*Meat+3*cheese.

New price of meat = $5

New budget line BL2

24=5*Meat+3*Cheese

New X axis intercept= 24/5= 4.8

New Y axis intercept= 24/3=8

Intercepts are (4.8,0) &( 0,8).

Assuming meat is a normal good, when price of meat falls, consumer demands more meat. So the demand curve is downward sloping.